OpenStudy (moongazer):

Find the equation of the line through the point (1, -5) and || to the x-axis.

OpenStudy (anonymous):

y+5=0

OpenStudy (moongazer):

how did you do it?

OpenStudy (moongazer):

Please explain.

OpenStudy (precal):

It goes through the value of -5 and since it needs to be parallel that means that the line has to be a horizontal line and the only way to write a horizontal line is y= number in this case it is y=-5 or y+5=0 (same line)

OpenStudy (anonymous):

parallel to x axis means constant value of y. so, y=-5 y+5=0

OpenStudy (anonymous):

its parallel to x axis u can say the equation is y=k putting the value of y k=-5 so,the equation is y=-5

OpenStudy (precal):

Sometimes it is easy to use graph paper and actually see what the point looks like and then draw in the line. Eventually you would have seen that you needed a horizontal line.